OpenStudy (anonymous):

z-6.2z-6=5.5(z+2.8) Whats z?

OpenStudy (aravindg):

-2

OpenStudy (aravindg):

giv medal to me

OpenStudy (aravindg):

:)

OpenStudy (anonymous):

-2

OpenStudy (anonymous):

ill give jimmy another chance

OpenStudy (anonymous):

z - 6.2z - 6 = 5.5z + 15.4 z - 6.2z - 5.5z = 15.4 + 6 -10.7z = 21.4 z = 21.4 / -10.7 z = -2

OpenStudy (anonymous):

ty dog!